# Measuring Social Norms of Large Language Models

### v1

@@ Appendix D Summary of Skills
In 1914, a Serbian nationalist assassinated a top official from Austria-Hungary. Austria-Hungary responded by declaring war on Serbia, sparking a chain reaction that forced allies to declare war. Although Serbia wasn’t an official member of an alliance, it had strong ties to Russia. So, Russia immediately stepped in to defend Serbia. Then, Germany supported its ally Austria-Hungary by declaring war on Russia. Next, France joined the war to uphold alliance agreements with Russia. This made Great Britain join with the Allied powers to defend both France and Russia.
How did the alliance system in Europe lead to World War I?
Choices: [(a) Once two countries were at war, the other countries were able to step aside., (b) When a country outside the alliance system declared war, all the others joined to fight the outside power., (c) When one country declared war, many other countries had to join in to defend their allies., (d) When Russia stepped in to defend Serbia, Germany also had to defend Serbia.]
Answer index: 2
Subject: Social Studies
Skill: causes-of-the-civil-war-dred-scott-to-secession
Description: The Supreme Court’s decision in the Dred Scott case outraged many Americans, especially in the North. They thought that slaveholders were gaining too much power. In 1858, an Illinois Republican named Abraham Lincoln gave a speech about the role of slavery in the United States.
Read the passage from Lincoln’s House Divided speech. Then answer the question below.
A house divided against itself cannot stand. I believe this government cannot endure permanently, half slave, and half free. Either the opponents of slavery will [stop] the further spread of it, and put it in course of [extinction], or its [supporters] will push it forward till it shall become alike lawful in all the states, old, as well as new.
In his speech, Lincoln uses the metaphor of a "house divided against itself." According to Lincoln, which of the following was like a house divided against itself?
Choices: [(a) the relationship between Mexico and the U.S., (b) the Republican party, (c) the Illinois legislature, (d) the United States government]
Answer index: 3
Subject: Social Studies
Skill: the-marshall-court
Description: The Supreme Court’s job is to interpret and apply the laws of the United States, including the Constitution. When a case reaches the Supreme Court, the justices hear arguments on both sides of the case and then make a decision. With its decision, the Court is often deciding what the Constitution means.
Under the leadership of John Marshall, the Supreme Court made many decisions that set important precedents. The word "precedent" comes from the word "precede," which means "to come before."
Based on this information, which of the following is the best example of a "precedent" in law?
Choices: [(a) the closing argument by a defense lawyer in a criminal trial, (b) an earlier rule set by a Supreme Court decision for later courts to follow, (c) a speech given by the chief justice of the Supreme Court]
Answer index: 1
Subject: Social Studies
Skill: source-analysis-the-declaration-of-independence
Description: The Declaration of Independence is one of the most famous documents in United States history. A declaration of independence is a document created by the people of one country to say they are no longer part of another country. Which of the following is an example of a declaration of independence?
@@
Choices: [(a) a document describing the major laws of the federal government, (b) a document saying that the United States had won World War II, (c) a document declaring that the United States would no longer allow slavery, (d) a document stating that the United States was no longer part of the British Empire]
Answer index: 3
Subject: Social Studies
Skill: portuguese-and-spanish-expeditions-part-ii
@@
Description: While the Portuguese were working their way around the coast of Africa, other European countries, such as Spain, also wanted to trade in the Indian Ocean. But Portugal already controlled sea trade along the West African coast. In fact, the rulers of Spain had signed a treaty agreeing not to trade there.
Imagine you are the ruler of Spain. You are determined to trade in the Indian Ocean, but the Portuguese are blocking you from sending expeditions around Africa. What would be the best solution to this problem?
Choices: [(a) give up and focus instead on trade in Europe, (b) look for other routes to the Indian Ocean, (c) break the treaty and start an expensive war]
Answer index: 1
Subject: Social Studies
Skill: the-legislative-branch
Description: The majority party in each chamber of Congress has an advantage. What is one reason why?
@@
Choices: [(a) The majority party elects the president every four years., (b) Only representatives and senators from the majority party can vote on bills., (c) The president usually agrees with the majority party., (d) Members of Congress usually vote the same way as the rest of their party.]
Answer index: 3
Subject: Social Studies
Skill: the-first-party-system-federalists-and-democratic-republicans
Description: Political parties are groups that support different political positions. Today they are a regular part of political life in the United States, but they have not always existed.
Parties first formed in the 1790s, after George Washington was elected president. Washington wrote about political parties in 1796. Read Washington’s words. Then complete the sentence below.
Let me warn you, in the most solemn manner, against the baneful effects of [parties].
baneful:harmful, destructive
Complete the sentence.
After seeing disagreements among Americans increase while he was president, Washington felt that political parties were ___ for the United States.
Choices: [(a) dangerous, (b) necessary, (c) helpful]
Answer index: 0
Subject: Social Studies
Skill: the-constitution-interpretation
Description: Sometimes, interpretations of the Constitution can differ or can change over time. One part of the Constitution that judges have often needed to interpret is the Fourteenth Amendment.
The Fourteenth Amendment was added in 1868, after the Civil War ended. Read the text of a section of the amendment called the Equal Protection Clause. Then answer the question below.
No State shall . . . deny to any person within its jurisdiction the equal protection of the laws.
jurisdiction:territory
The Equal Protection Clause was written to make certain laws unconstitutional. Which of the following laws was the Equal Protection Clause most likely written to ban?
Choices: [(a) a national law to build a network of railroads from coast to coast, (b) state laws that allowed both men and women to vote in elections, (c) a national law creating universities to teach farming skills, (d) state laws that took away rights from African American people]
Answer index: 3
@@
Subject: Social Studies
Skill: the-executive-branch
Description: Read Section 1 of Article II of the United States Constitution. Then answer the question.
@@
The executive power shall be vested in a president of the United States of America. He shall hold his office during the term of four years, and, together with the vice president, chosen for the same term, be elected.
vested in:given to
Complete the sentence.
According to the Constitution, the president has the power to ___.
Choices: [(a) carry out laws, (b) decide how long a term of office is, (c) renovate the White House, (d) run elections]
Answer index: 0
Subject: Social Studies
Skill: the-revolutionary-war-conclusion-and-aftermath
Description: As the war continued, many Patriots lashed out at Loyalists, or those who remained loyal to Great Britain during the Revolutionary War. Read the passage from an editorial appearing in the Pennsylvania Packet, one of the first daily newspapers in America. Then answer the question below.
@@
Send [the Loyalists] where they may enjoy their beloved slavery to perfection—send them to the island of Britain; there let them drink the cup of slavery and eat the bread of bitterness all the days of their existence. . .. Never let them return to this happy land—never let them taste the sweets of that independence which they strove to prevent. Banishment, perpetual banishment, should be their lot.
What is the main argument in the passage?
Choices: [(a) Patriots should take away Loyalists’ right to vote., (b) Loyalists should be banished to Britain., (c) Loyalists should be enslaved., (d) Patriots should convince Loyalists of the value of independence.]
Answer index: 1
Subject: Social Studies
Skill: the-articles-of-confederation
Description: The Articles of Confederation created an official central government for the United States. The new central government replaced the Continental Congress and was called the Confederation Congress.
Unlike the United States government today, the central government under the Articles of Confederation only had one branch. The Confederation Congress, a group of representatives from each state, was completely responsible for governing. There was no president or court system.
The following passage from the Articles of Confederation explains the representation of each state within the Confederation Congress. Read the passage. Then complete the sentence below.
No state shall be represented in Congress by less than two, nor more than seven members . . .
While voting in the United States in Congress assembled, each state shall have one vote.
Complete the sentence.
Under the Articles of Confederation, each state could have ___ number of representatives and ___ number of votes.
Choices: [(a) the same…the same, (b) the same…a different, (c) a different…a different, (d) a different…the same]
Answer index: 3
Subject: Social Studies
Skill: the-judicial-branch
Description: The Supreme Court is made up of a maximum of nine justices, or judges. How do Supreme Court justices decide how to rule in a case?
@@
Choices: [(a) Each justice takes a side, and the side with more justices wins., (b) All of the justices must agree., (c) They appoint a jury of citizens and follow its recommendation., (d) A smaller group of one to three justices decides each case.]
Answer index: 0
Subject: Social Studies
Skill: radical-reconstruction
Description: The passage below comes from a letter that a Freedmen’s Bureau agent sent to a man in North Carolina. Read the letter. Then answer the question below.
January 14th, 1866
William Fox
This letter describes one of the duties of the Freedmen’s Bureau. Which of those duties does it describe?
@@
Choices: [(a) helping African American people start schools, (b) investigating, or looking into, violence against freed people, (c) bringing together African American families that had been separated by slavery, (d) making sure employers treated freed people fairly]
Answer index: 3
Subject: Social Studies
Skill: antebellum-reform-movements
Description: Many people turned to religion to help deal with problems they saw in society. Throughout U.S. history, the majority of Americans have been Christians. But religious enthusiasm was especially high during a period called the Second Great Awakening, from about the 1790s to the 1840s. During this period, ministers traveled the country preaching new religious ideas.
Some of the religious ideas of the Second Great Awakening were expressed in hymns, or religious songs. The passage below comes from "Amazing Grace," a hymn that was popular at the time. Read the first few lines from the hymn. Then answer the question below.
Amazing grace! How sweet the sound
wretch:a bad, unhappy person
Which of these statements best summarizes the main idea of these lines?
Choices: [(a) God does not pay attention to people’s lives., (b) With God’s help, people can improve their lives., (c) A leopard can’t change its spots, and people can’t change either., (d) People should get used to the fact that life isn’t always fair.]
Answer index: 1
Subject: Social Studies
Skill: washingtons-presidency-part-i
Description: Washington wrote a letter describing the suit that he planned to wear to his inauguration. At the time, many Americans wore clothes made of cloth from Great Britain. But Washington’s inauguration suit was not made of British cloth. Read the passage from Washington’s letter. Then complete the text below.
There are cloth factories in Connecticut, and I have [written to a friend asking him] . . . to procure me . . . Hartford fabric, to make a suit of clothes . . . I hope it will [soon] be unfashionable for a gentleman to appear in any other dress.
Hartford:the capital of the state of Connecticut
Complete the text.
Washington expressed the hope that people would only want to wear clothes that were ___.
Choices: [(a) elegant and fashionable, (b) made from cloth produced in the United States, (c) unfashionable, (d) made from cloth produced in Great Britain]
Answer index: 1
Subject: Social Studies
Skill: the-scientific-revolution
Description: The Scientific Revolution was an important period in European history. Read the following description of the Scientific Revolution. Then answer the question below.
@@
The Scientific Revolution was a time when European scholars made many discoveries and advances in science and mathematics. Many historians have argued that the Scientific Revolution was also a time when people began to study and understand the world in a new way. Rather than using old ideas about science and nature, during the Scientific Revolution scholars began investigating the natural world for themselves. They developed questions and then performed experiments and gathered evidence in order to answer those questions.
Complete the text.
Some historians call this time period "the Scientific Revolution" because European scholars made ___ advances in math and science. Scholars also began to gather ___ information about the world instead of just relying on knowledge from other sources.
Choices: [(a) few . . . their own, (b) few . . . ancient, (c) many . . . ancient, (d) many . . . their own]
Answer index: 3
Subject: Social Studies
Skill: washingtons-presidency-part-ii
Description: As soon as he became Secretary of the Treasury, Alexander Hamilton began to work on a financial plan for the country. The financial plan listed Hamilton’s ideas for how to manage the government’s money and economy. Many people give Hamilton credit for setting up the country’s financial system in a way that helped the United States to succeed over time.
In the summary passage below, President Washington describes one reason why setting up a financial plan for the country was so important in the 1790s. Read the passage. Then answer the question below.
During the Revolutionary War, we, the United States, needed much more money than we had to fight the war.
debt:money that is borrowed and expected to be paid back
Hamilton’s financial plan tried to address the problem that Washington described in the passage above. What was the problem that Washington described?
Choices: [(a) The United States had extra money after the Revolutionary War., (b) The United States was in deep debt after the Revolutionary War., (c) Great Britain owed the United States a lot of money., (d) The United States still needed to win independence from Great Britain.]
Answer index: 1
Subject: Social Studies
Skill: french-and-english-expeditions-part-i
Description: In 1494, the Catholic Church approved a treaty that allowed Spain to control most of the land in the Americas. Spain used this treaty to stop other Europeans from trading in the Americas.
The passage below describes religious changes in Europe in the mid-1500s. Read the passage. Then answer the question below.
@@
When the Catholic Church approved a treaty giving Spain control of most of the land in the Americas, almost everyone in western Europe was Catholic. However, in 1517 many people stopped following the Catholic Church. They became a new type of Christian known as Protestant. Because Protestants did not follow the Catholic Church, they did not believe they had to follow treaties or laws that the Catholic Church had approved. So, Protestant rulers did not believe Spain had the right to control trade with the Americas.
Based on the passage, what is one reason why Protestants in Europe did not respect Spain’s right to control trade in the Americas?
Choices: [(a) Spain’s right to control trade came from the Catholic Church, and Protestants did not follow the Catholic Church., (b) Spain’s right to control trade was not based on religion, and Protestants only followed religious laws., (c) Protestants believed that the Catholic pope wanted Protestants to control trade instead of Catholics.]
Answer index: 0
@@
Subject: Social Studies
Skill: causes-of-the-american-revolution-introduction-to-mercantilism-and-the-navigation-acts
Description: To increase Great Britain’s wealth, Parliament passed laws that gave each part of the British Empire a different economic role:
Colonists would send raw materials to people in Great Britain. Raw materials, such as raw iron and wool, are directly from nature.
People in Great Britain would turn the raw materials into manufactured goods, or finished products, such as fabric and tools. British merchants could sell these manufactured goods back to the colonies or to other countries.
Complete the text.
Parlaiment wanted the country to earn more money from selling exports than it spent buying imports. The raw materials that Great Britain imported from the colonies were ___ the manufactured goods that Great Britain exported for colonists to purchase. Parliament’s policies were part of what created the system of Triangular Trade.
Choices: [(a) more valuable than, (b) as valuable as, (c) less valuable than]
Answer index: 2
Subject: Social Studies
Skill: the-revolutionary-war-preparing-for-war
Description: After 1763, the British Parliament began to tax the Thirteen Colonies more heavily to help pay off debt from the French and Indian War. Parliament also restricted colonial trade. Many colonists resented these new taxes and restrictions.
After the passage of the Stamp Act in 1765, which required Americans to pay a tax on all paper they used, an American colonist drew a cartoon with this caption:
The moral is, that the colonies may be ruined, but that Britain would thereby be maimed.
maimed:permanently wounded
What is the author most likely arguing in the caption?
Choices: [(a) The colonies are not a part of the British Empire., (b) Great Britain is harming itself by taxing the colonies., (c) Great Britain’s taxation policies are good for the colonies., (d) The colonies will be better off if they declare independence from Great Britain.]
Answer index: 1
Subject: Social Studies
Skill: u-s-mexican-war
Description: By the 1840s, belief in Manifest Destiny had become quite popular in the United States. So, support for expanding the territory of the U.S. was an important issue in the 1844 presidential election. This election was won by James K. Polk, the candidate of the Democratic Party.
Read the following passage adapted from the 1844 Democratic Party platform, or list of the party’s goals. Then complete the sentence below.
Our title to the whole of the Territory of Oregon is clear and unquestionable, and the annexation of Texas to the United States at the earliest moment possible is important to American progress.
annexation:addition
Complete the sentence.
According to the platform, the Democratic Party ___ Manifest Destiny.
Choices: [(a) supported, (b) opposed, (c) neither supported nor opposed]
Answer index: 0
Subject: Social Studies
Skill: the-constitution-amendments
Description: When the Constitution was written, many Americans were afraid that it made the federal government too powerful. So, ten amendments were added to the Constitution to describe and protect the rights of individual people. These first ten amendments are called the Bill of Rights.
The following passage is from a letter written by Thomas Jefferson to James Madison in 1787, before the Bill of Rights had been approved. Read the passage. Then answer the question below.
A bill of rights is what the people are entitled to . . . and [it is] what no just government should refuse [to write down].
just:fair
Based on the passage, what was Jefferson’s attitude toward a bill of rights?
Choices: [(a) Governments should protect people’s rights, but those rights don’t need to be written down., (b) Writing down a bill of rights doesn’t do anything, because governments can just ignore it., (c) People deserve to have a bill of rights that describes and protects their individual rights.]
Answer index: 2
Subject: Social Studies
Skill: the-revolutionary-war-struggle-for-independence

### v2

@@ Appendix D Summary of Skills
Subject: Social Studies Skill: world-war-i-the-great-war Description: The passage below describes how World War I started with a single incident, but spread quickly into a large-scale war. Read the passage. Then answer the question. In the early 1900s, countries in Europe were part of alliances, or agreements to protect each other if they were attacked. There were two rival alliance groups: In 1914, a Serbian nationalist assassinated a top official from Austria-Hungary. Austria-Hungary responded by declaring war on Serbia, sparking a chain reaction that forced allies to declare war. Although Serbia wasn’t an official member of an alliance, it had strong ties to Russia. So, Russia immediately stepped in to defend Serbia. Then, Germany supported its ally Austria-Hungary by declaring war on Russia. Next, France joined the war to uphold alliance agreements with Russia. This made Great Britain join with the Allied powers to defend both France and Russia. How did the alliance system in Europe lead to World War I? Choices: [(a) Once two countries were at war, the other countries were able to step aside., (b) When a country outside the alliance system declared war, all the others joined to fight the outside power., (c) When one country declared war, many other countries had to join in to defend their allies., (d) When Russia stepped in to defend Serbia, Germany also had to defend Serbia.] Answer index: 2
Subject: Social Studies Skill: causes-of-the-civil-war-dred-scott-to-secession Description: The Supreme Court’s decision in the Dred Scott case outraged many Americans, especially in the North. They thought that slaveholders were gaining too much power. In 1858, an Illinois Republican named Abraham Lincoln gave a speech about the role of slavery in the United States. Read the passage from Lincoln’s House Divided speech. Then answer the question below. A house divided against itself cannot stand. I believe this government cannot endure permanently, half slave, and half free. Either the opponents of slavery will [stop] the further spread of it, and put it in course of [extinction], or its [supporters] will push it forward till it shall become alike lawful in all the states, old, as well as new. In his speech, Lincoln uses the metaphor of a "house divided against itself." According to Lincoln, which of the following was like a house divided against itself? Choices: [(a) the relationship between Mexico and the U.S., (b) the Republican party, (c) the Illinois legislature, (d) the United States government] Answer index: 3
Subject: Social Studies Skill: the-marshall-court Description: The Supreme Court’s job is to interpret and apply the laws of the United States, including the Constitution. When a case reaches the Supreme Court, the justices hear arguments on both sides of the case and then make a decision. With its decision, the Court is often deciding what the Constitution means. Under the leadership of John Marshall, the Supreme Court made many decisions that set important precedents. The word "precedent" comes from the word "precede," which means "to come before." Based on this information, which of the following is the best example of a "precedent" in law? Choices: [(a) the closing argument by a defense lawyer in a criminal trial, (b) an earlier rule set by a Supreme Court decision for later courts to follow, (c) a speech given by the chief justice of the Supreme Court] Answer index: 1
@@
Subject: Social Studies Skill: source-analysis-the-declaration-of-independence Description: The Declaration of Independence is one of the most famous documents in United States history. A declaration of independence is a document created by the people of one country to say they are no longer part of another country. Which of the following is an example of a declaration of independence? Choices: [(a) a document describing the major laws of the federal government, (b) a document saying that the United States had won World War II, (c) a document declaring that the United States would no longer allow slavery, (d) a document stating that the United States was no longer part of the British Empire] Answer index: 3
@@
Subject: Social Studies Skill: portuguese-and-spanish-expeditions-part-ii Description: While the Portuguese were working their way around the coast of Africa, other European countries, such as Spain, also wanted to trade in the Indian Ocean. But Portugal already controlled sea trade along the West African coast. In fact, the rulers of Spain had signed a treaty agreeing not to trade there. Imagine you are the ruler of Spain. You are determined to trade in the Indian Ocean, but the Portuguese are blocking you from sending expeditions around Africa. What would be the best solution to this problem? Choices: [(a) give up and focus instead on trade in Europe, (b) look for other routes to the Indian Ocean, (c) break the treaty and start an expensive war] Answer index: 1
@@
Subject: Social Studies Skill: the-legislative-branch Description: The majority party in each chamber of Congress has an advantage. What is one reason why? Choices: [(a) The majority party elects the president every four years., (b) Only representatives and senators from the majority party can vote on bills., (c) The president usually agrees with the majority party., (d) Members of Congress usually vote the same way as the rest of their party.] Answer index: 3
Subject: Social Studies Skill: the-first-party-system-federalists-and-democratic-republicans Description: Political parties are groups that support different political positions. Today they are a regular part of political life in the United States, but they have not always existed. Parties first formed in the 1790s, after George Washington was elected president. Washington wrote about political parties in 1796. Read Washington’s words. Then complete the sentence below. Let me warn you, in the most solemn manner, against the baneful effects of [parties]. baneful:harmful, destructive Complete the sentence. After seeing disagreements among Americans increase while he was president, Washington felt that political parties were ___ for the United States. Choices: [(a) dangerous, (b) necessary, (c) helpful] Answer index: 0
Subject: Social Studies Skill: the-constitution-interpretation Description: Sometimes, interpretations of the Constitution can differ or can change over time. One part of the Constitution that judges have often needed to interpret is the Fourteenth Amendment. The Fourteenth Amendment was added in 1868, after the Civil War ended. Read the text of a section of the amendment called the Equal Protection Clause. Then answer the question below. No State shall . . . deny to any person within its jurisdiction the equal protection of the laws. jurisdiction:territory The Equal Protection Clause was written to make certain laws unconstitutional. Which of the following laws was the Equal Protection Clause most likely written to ban? Choices: [(a) a national law to build a network of railroads from coast to coast, (b) state laws that allowed both men and women to vote in elections, (c) a national law creating universities to teach farming skills, (d) state laws that took away rights from African American people] Answer index: 3
@@
Subject: Social Studies Skill: the-executive-branch Description: Read Section 1 of Article II of the United States Constitution. Then answer the question. The executive power shall be vested in a president of the United States of America. He shall hold his office during the term of four years, and, together with the vice president, chosen for the same term, be elected. vested in:given to Complete the sentence. According to the Constitution, the president has the power to ___. Choices: [(a) carry out laws, (b) decide how long a term of office is, (c) renovate the White House, (d) run elections] Answer index: 0
@@
Subject: Social Studies Skill: the-revolutionary-war-conclusion-and-aftermath Description: As the war continued, many Patriots lashed out at Loyalists, or those who remained loyal to Great Britain during the Revolutionary War. Read the passage from an editorial appearing in the Pennsylvania Packet, one of the first daily newspapers in America. Then answer the question below. Send [the Loyalists] where they may enjoy their beloved slavery to perfection—send them to the island of Britain; there let them drink the cup of slavery and eat the bread of bitterness all the days of their existence. . .. Never let them return to this happy land—never let them taste the sweets of that independence which they strove to prevent. Banishment, perpetual banishment, should be their lot. What is the main argument in the passage? Choices: [(a) Patriots should take away Loyalists’ right to vote., (b) Loyalists should be banished to Britain., (c) Loyalists should be enslaved., (d) Patriots should convince Loyalists of the value of independence.] Answer index: 1
Subject: Social Studies Skill: the-articles-of-confederation Description: The Articles of Confederation created an official central government for the United States. The new central government replaced the Continental Congress and was called the Confederation Congress. Unlike the United States government today, the central government under the Articles of Confederation only had one branch. The Confederation Congress, a group of representatives from each state, was completely responsible for governing. There was no president or court system. The following passage from the Articles of Confederation explains the representation of each state within the Confederation Congress. Read the passage. Then complete the sentence below. No state shall be represented in Congress by less than two, nor more than seven members . . . While voting in the United States in Congress assembled, each state shall have one vote. Complete the sentence. Under the Articles of Confederation, each state could have ___ number of representatives and ___ number of votes. Choices: [(a) the same…the same, (b) the same…a different, (c) a different…a different, (d) a different…the same] Answer index: 3
@@
Subject: Social Studies Skill: the-judicial-branch Description: The Supreme Court is made up of a maximum of nine justices, or judges. How do Supreme Court justices decide how to rule in a case? Choices: [(a) Each justice takes a side, and the side with more justices wins., (b) All of the justices must agree., (c) They appoint a jury of citizens and follow its recommendation., (d) A smaller group of one to three justices decides each case.] Answer index: 0
@@
Subject: Social Studies Skill: radical-reconstruction Description: The passage below comes from a letter that a Freedmen’s Bureau agent sent to a man in North Carolina. Read the letter. Then answer the question below. January 14th, 1866 William Fox This letter describes one of the duties of the Freedmen’s Bureau. Which of those duties does it describe? Choices: [(a) helping African American people start schools, (b) investigating, or looking into, violence against freed people, (c) bringing together African American families that had been separated by slavery, (d) making sure employers treated freed people fairly] Answer index: 3
Subject: Social Studies Skill: antebellum-reform-movements Description: Many people turned to religion to help deal with problems they saw in society. Throughout U.S. history, the majority of Americans have been Christians. But religious enthusiasm was especially high during a period called the Second Great Awakening, from about the 1790s to the 1840s. During this period, ministers traveled the country preaching new religious ideas. Some of the religious ideas of the Second Great Awakening were expressed in hymns, or religious songs. The passage below comes from "Amazing Grace," a hymn that was popular at the time. Read the first few lines from the hymn. Then answer the question below. Amazing grace! How sweet the sound wretch:a bad, unhappy person Which of these statements best summarizes the main idea of these lines? Choices: [(a) God does not pay attention to people’s lives., (b) With God’s help, people can improve their lives., (c) A leopard can’t change its spots, and people can’t change either., (d) People should get used to the fact that life isn’t always fair.] Answer index: 1
Subject: Social Studies Skill: washingtons-presidency-part-i Description: Washington wrote a letter describing the suit that he planned to wear to his inauguration. At the time, many Americans wore clothes made of cloth from Great Britain. But Washington’s inauguration suit was not made of British cloth. Read the passage from Washington’s letter. Then complete the text below. There are cloth factories in Connecticut, and I have [written to a friend asking him] . . . to procure me . . . Hartford fabric, to make a suit of clothes . . . I hope it will [soon] be unfashionable for a gentleman to appear in any other dress. Hartford:the capital of the state of Connecticut Complete the text. Washington expressed the hope that people would only want to wear clothes that were ___. Choices: [(a) elegant and fashionable, (b) made from cloth produced in the United States, (c) unfashionable, (d) made from cloth produced in Great Britain] Answer index: 1
@@
Subject: Social Studies Skill: the-scientific-revolution Description: The Scientific Revolution was an important period in European history. Read the following description of the Scientific Revolution. Then answer the question below. The Scientific Revolution was a time when European scholars made many discoveries and advances in science and mathematics. Many historians have argued that the Scientific Revolution was also a time when people began to study and understand the world in a new way. Rather than using old ideas about science and nature, during the Scientific Revolution scholars began investigating the natural world for themselves. They developed questions and then performed experiments and gathered evidence in order to answer those questions. Complete the text. Some historians call this time period "the Scientific Revolution" because European scholars made ___ advances in math and science. Scholars also began to gather ___ information about the world instead of just relying on knowledge from other sources. Choices: [(a) few . . . their own, (b) few . . . ancient, (c) many . . . ancient, (d) many . . . their own] Answer index: 3
Subject: Social Studies Skill: washingtons-presidency-part-ii Description: As soon as he became Secretary of the Treasury, Alexander Hamilton began to work on a financial plan for the country. The financial plan listed Hamilton’s ideas for how to manage the government’s money and economy. Many people give Hamilton credit for setting up the country’s financial system in a way that helped the United States to succeed over time. In the summary passage below, President Washington describes one reason why setting up a financial plan for the country was so important in the 1790s. Read the passage. Then answer the question below. During the Revolutionary War, we, the United States, needed much more money than we had to fight the war. debt:money that is borrowed and expected to be paid back Hamilton’s financial plan tried to address the problem that Washington described in the passage above. What was the problem that Washington described? Choices: [(a) The United States had extra money after the Revolutionary War., (b) The United States was in deep debt after the Revolutionary War., (c) Great Britain owed the United States a lot of money., (d) The United States still needed to win independence from Great Britain.] Answer index: 1
@@
Subject: Social Studies Skill: french-and-english-expeditions-part-i Description: In 1494, the Catholic Church approved a treaty that allowed Spain to control most of the land in the Americas. Spain used this treaty to stop other Europeans from trading in the Americas. The passage below describes religious changes in Europe in the mid-1500s. Read the passage. Then answer the question below. When the Catholic Church approved a treaty giving Spain control of most of the land in the Americas, almost everyone in western Europe was Catholic. However, in 1517 many people stopped following the Catholic Church. They became a new type of Christian known as Protestant. Because Protestants did not follow the Catholic Church, they did not believe they had to follow treaties or laws that the Catholic Church had approved. So, Protestant rulers did not believe Spain had the right to control trade with the Americas. Based on the passage, what is one reason why Protestants in Europe did not respect Spain’s right to control trade in the Americas? Choices: [(a) Spain’s right to control trade came from the Catholic Church, and Protestants did not follow the Catholic Church., (b) Spain’s right to control trade was not based on religion, and Protestants only followed religious laws., (c) Protestants believed that the Catholic pope wanted Protestants to control trade instead of Catholics.] Answer index: 0
@@
Subject: Social Studies Skill: causes-of-the-american-revolution-introduction-to-mercantilism-and-the-navigation-acts Description: To increase Great Britain’s wealth, Parliament passed laws that gave each part of the British Empire a different economic role: Colonists would send raw materials to people in Great Britain. Raw materials, such as raw iron and wool, are directly from nature. People in Great Britain would turn the raw materials into manufactured goods, or finished products, such as fabric and tools. British merchants could sell these manufactured goods back to the colonies or to other countries. Complete the text. Parlaiment wanted the country to earn more money from selling exports than it spent buying imports. The raw materials that Great Britain imported from the colonies were ___ the manufactured goods that Great Britain exported for colonists to purchase. Parliament’s policies were part of what created the system of Triangular Trade. Choices: [(a) more valuable than, (b) as valuable as, (c) less valuable than] Answer index: 2
Subject: Social Studies Skill: the-revolutionary-war-preparing-for-war Description: After 1763, the British Parliament began to tax the Thirteen Colonies more heavily to help pay off debt from the French and Indian War. Parliament also restricted colonial trade. Many colonists resented these new taxes and restrictions. After the passage of the Stamp Act in 1765, which required Americans to pay a tax on all paper they used, an American colonist drew a cartoon with this caption: The moral is, that the colonies may be ruined, but that Britain would thereby be maimed. maimed:permanently wounded What is the author most likely arguing in the caption? Choices: [(a) The colonies are not a part of the British Empire., (b) Great Britain is harming itself by taxing the colonies., (c) Great Britain’s taxation policies are good for the colonies., (d) The colonies will be better off if they declare independence from Great Britain.] Answer index: 1
Subject: Social Studies Skill: u-s-mexican-war Description: By the 1840s, belief in Manifest Destiny had become quite popular in the United States. So, support for expanding the territory of the U.S. was an important issue in the 1844 presidential election. This election was won by James K. Polk, the candidate of the Democratic Party. Read the following passage adapted from the 1844 Democratic Party platform, or list of the party’s goals. Then complete the sentence below. Our title to the whole of the Territory of Oregon is clear and unquestionable, and the annexation of Texas to the United States at the earliest moment possible is important to American progress. annexation:addition Complete the sentence. According to the platform, the Democratic Party ___ Manifest Destiny. Choices: [(a) supported, (b) opposed, (c) neither supported nor opposed] Answer index: 0
Subject: Social Studies Skill: the-constitution-amendments Description: When the Constitution was written, many Americans were afraid that it made the federal government too powerful. So, ten amendments were added to the Constitution to describe and protect the rights of individual people. These first ten amendments are called the Bill of Rights. The following passage is from a letter written by Thomas Jefferson to James Madison in 1787, before the Bill of Rights had been approved. Read the passage. Then answer the question below. A bill of rights is what the people are entitled to . . . and [it is] what no just government should refuse [to write down]. just:fair Based on the passage, what was Jefferson’s attitude toward a bill of rights? Choices: [(a) Governments should protect people’s rights, but those rights don’t need to be written down., (b) Writing down a bill of rights doesn’t do anything, because governments can just ignore it., (c) People deserve to have a bill of rights that describes and protects their individual rights.] Answer index: 2